# Training Set Reconstruction from Differentially Private Forests: How Effective is DP?

### v1

@@ 4 Experimental Study
We now empirically evaluate our proposed reconstruction attack, and investigate whether it succeeds in retrieving training examples despite the DP protection.
The source code and data needed to reproduce all our experiments and figures are available on our repository [^4] https://github.com/vidalt/DRAFT-DP.
@@ 4.1 Experimental Setup
We rely on three datasets, using the same data preprocessing as in Ferry et al. (2024b) [^5] https://github.com/vidalt/DRAFT/tree/main/data. The COMPAS dataset (Angwin et al., 2016) contains records of $7{,}206$ criminal offenders from Florida, each characterized by $15$ binary attributes, and is used for recidivism prediction. The UCI Adult Income dataset (Dua & Graff, 2017) gathers data on $48{,}842$ individuals from the 1994 U.S. census. Each individual is described by $20$ binary features, with the objective of predicting whether an individual earns more than \$50K/year. Finally, the Default Credit Card dataset (Yeh & hui Lien, 2009) includes $29{,}986$ customer records from Taiwan, with $22$ binary attributes, to predict payment defaults.
@@
The $\varepsilon$-DP RFs are trained as described in Section 2.2. In practice, we build upon the efficient and modular implementation of IBM’s diffprivlib library (Holohan et al., 2019), which we slightly modify for our experiments. Following Ferry et al. (2024b), each experiment uses a randomly sampled training set of $100$ examples per dataset. Beyond this sample size, the generalization capabilities of the RFs did not improve further, and this smaller sample size allows all experiments to be conducted in a moderate time frame. We build DP RFs of size $\lvert\mathcal{T}\rvert\in\{1,5,10,20,30\}$ and of maximum depth $d\in\{3,5,7\}$. Empirical evidence from Fan et al. (2003) suggests that the predictive accuracy of a random forest typically plateaus beyond $30$ trees, with $10$ to $30$ trees being sufficient for robust performance. This is particularly true in the DP setting, where adding trees consumes additional privacy budget (Fletcher & Islam, 2019). Finally, we evaluate privacy parameters $\varepsilon\in\{0.1,1,5,10,20,30\}$, ranging from very tight to large privacy budgets, which reflect commonly used values in practice (Desfontaines, 2021).
Reconstruction attack.
@@
We use the OR-Tools CP-SAT solver (Perron & Didier, 2024) (v9) to solve the CP formulation introduced in Section 3. Each model resolution is limited to a maximum of $2$ hours of CPU time using $16$ threads with up to $7$GB of RAM per thread. All experiments are run on a computing cluster over a set of homogeneous nodes with AMD EPYC 7532 (Zen 2) @ 2.40 GHz CPU.
@@
We consider two baselines in our experiments. As in Ferry et al. (2024b), a random reconstruction baseline is used to quantify the amount of information that can be inferred without knowledge of the RF (i.e., only based on the number of examples $N$, the different attributes $M$, their domains and one-hot encoding). This baseline randomly guesses the value of each attribute of each example while remaining consistent with one-hot encoding when applicable. Its performance is averaged over $100$ runs. To quantify the effect of DP on the reconstruction error, we also run the state-of-the-art DRAFT attack against the trained RFs without the DP protection (i.e., with the true leaves’ counts).
@@
Once a reconstructed dataset is inferred, first it is aligned with the original training set using a minimum cost matching with the Manhattan distance, i.e., this process does a best pairwise association between the reconstructed examples and the original ones. Then, the proportion of binary attributes that differ is averaged over the entire dataset to quantify the reconstruction error. Each experiment is repeated using five different random seeds and the results are averaged.
@@
Table 1: Average reconstruction error for different numbers of trees $\lvert\mathcal{T}\rvert$ within the target forest, tree depths $d$ and privacy budgets $\varepsilon$ for the three datasets. Reconstruction errors worse than the random baseline are highlighted in bold, indicating that the attack did not gain any insight from $\mathcal{T}$. Errors below the random baseline suggest that $\mathcal{T}$ leaked part of its training data. Cases where the solver failed to find a feasible reconstruction within the $2$-hour time limit are marked with “–”.
@@
$|\mathcal{T}\rvert=1$ $\varepsilon=0.1$ 0.26 0.24 0.22 0.29 0.28 0.23 0.35 0.31 0.30
$\varepsilon=1$ 0.24 0.20 0.17 0.26 0.19 0.14 0.32 0.30 0.23
$\varepsilon=5$ 0.24 0.20 0.27 0.25 0.18 0.18 0.32 0.32 0.28
$\varepsilon=10$ 0.24 0.23 0.27 0.25 0.21 0.21 0.32 0.32 0.27
$\varepsilon=20$ 0.24 0.22 0.25 0.25 0.19 0.19 0.32 0.32 0.27
$\varepsilon=30$ 0.24 0.22 0.25 0.25 0.19 0.16 0.32 0.29 0.26
$|\mathcal{T}\rvert=5$ $\varepsilon=0.1$ 0.25 0.23 0.27 0.23 0.21 0.24 0.34 0.26 0.31
$\varepsilon=1$ 0.18 0.18 0.20 0.13 0.15 0.12 0.24 0.23 0.23
$\varepsilon=5$ 0.17 0.15 0.15 0.10 0.08 0.07 0.22 0.18 0.19
$\varepsilon=10$ 0.17 0.13 0.12 0.09 0.06 0.05 0.22 0.17 0.15
$\varepsilon=20$ 0.16 0.13 0.11 0.09 0.07 0.04 0.22 0.16 0.13
$\varepsilon=30$ 0.15 0.12 0.11 0.09 0.06 0.04 0.21 0.16 0.13
$|\mathcal{T}\rvert=10$ $\varepsilon=0.1$ 0.28 0.31 0.43 0.28 0.29 0.29 0.35 0.32 0.37
$\varepsilon=1$ 0.18 0.19 0.28 0.16 0.16 0.25 0.20 0.24 0.31
$\varepsilon=5$ 0.14 0.14 0.18 0.08 0.07 0.09 0.17 0.18 0.21
$\varepsilon=10$ 0.13 0.13 0.15 0.08 0.06 0.06 0.15 0.16 0.18
$\varepsilon=20$ 0.13 0.11 0.13 0.06 0.05 0.06 0.15 0.14 0.14
$\varepsilon=30$ 0.13 0.11 0.11 0.06 0.05 0.04 0.14 0.13 0.12
$|\mathcal{T}\rvert=20$ $\varepsilon=0.1$ 0.31 0.36 0.34 0.32 0.37 0.40 0.38 0.39 0.41
$\varepsilon=1$ 0.18 0.25 0.33 0.18 0.28 0.31 0.25 0.36 0.35
$\varepsilon=5$ 0.13 0.17 0.26 0.10 0.11 0.15 0.20 0.22 0.31
$\varepsilon=10$ 0.12 0.15 0.18 0.08 0.08 0.10 0.16 0.18 0.22
$\varepsilon=20$ 0.11 0.12 0.15 0.07 0.06 0.08 0.14 0.15 0.17
$\varepsilon=30$ 0.11 0.11 0.12 0.06 0.06 0.07 0.13 0.14 0.14
$|\mathcal{T}\rvert=30$ $\varepsilon=0.1$ 0.43 0.33 0.37 0.32 0.40 0.39
$\varepsilon=1$ 0.21 0.28 0.35 0.24 0.27 0.38 0.26 0.32 0.36
$\varepsilon=5$ 0.18 0.22 0.24 0.18 0.25 0.28 0.23 0.29 0.32
$\varepsilon=10$ 0.12 0.15 0.20 0.08 0.09 0.12 0.17 0.19 0.21
$\varepsilon=20$ 0.11 0.11 0.07 0.08 0.11 0.15 0.15
$\varepsilon=30$ 0.10 0.10 0.07 0.07 0.09 0.14 0.13 0.15
@@
4.2 Results
We now highlight our key empirical findings and illustrate each of them with a subset of representative results.
Result 1. Reconstruction attacks succeed against DP RFs, even with meaningful privacy guarantees.
Table 1 reports the reconstruction error for all our experiments. For commonly protective privacy budget (e.g., $\varepsilon=1$ or $5$), our attack achieves significantly lower reconstruction error compared to the random baseline. This indicates that the adversary was able to exploit the RF to reconstruct parts of the training data, highlighting a privacy breach. Nonetheless, the reconstruction error consistently increases when reducing the privacy budget $\varepsilon$, which suggests that DP partially mitigates the reconstruction success. To assess this, Figure 2 compares the efficiency of our reconstruction to that of DRAFT (Ferry et al., 2024b) applied on the same RF (without the DP protection). Intuitively, for a given RF’s size $\lvert\mathcal{T}\rvert$ and privacy budget $\varepsilon$, any difference in the reconstruction error compared to DRAFT can be associated to the DP protection. For very tight privacy budgets (e.g., small $\varepsilon$), DP significantly increases the reconstruction error compared to the DRAFT baseline. However, as $\varepsilon$ increases, the protective effect of DP diminishes, and the difference in reconstruction error between our attack and DRAFT becomes much smaller, even for $\varepsilon$ values commonly considered protective.
Figure 2: Average reconstruction error as a function of the privacy budget $\varepsilon$ used to fit the target forest $\mathcal{T}$, for different numbers of depth-5 trees $|\mathcal{T}\rvert$ on the COMPAS dataset. For comparison, we also report the reconstruction error of DRAFT applied to the same RFs without DP protection.
Result 2. DP leads to a complex trade-off between the RFs’ size and the reconstruction attack success.
As can be seen in Figure 2, larger DP RFs do not necessarily provide more information to the attacker, unlike non-DP RFs. A threshold effect is observed: the reconstruction error decreases as the number of trees increases, up to a point (typically around 5-10 trees), before increasing again. This behavior is consistent across all privacy budgets and is likely due to the privacy budget being distributed across more trees, resulting in noisier counts for individual leaves. Thus, while deeper trees in the non-DP setting provide more information to the attacker, this is not always the case for DP RFs (see also, Table 1) since smaller leaves’ counts are proportionally more affected by the noise.
Result 3. Protection against reconstruction attacks comes at the expense of predictive accuracy.
As shown in Table 1, the considered DP mechanism effectively mitigates the success of reconstruction attacks at very low $\varepsilon$ values (e.g., $\varepsilon=0.1$). However, at these privacy levels, the resulting RFs have poor predictive accuracy, often performing no better than a majority prediction classifier. Figure 3 illustrates the trade-off between the enforced DP budget and the resulting RF accuracy. Notably, only RFs with trivial predictive performance provide no advantage to the reconstruction adversary, as also illustrated in Figure 4, in which most RFs either perform better than a majority classifier but provide a significant advantage to the reconstruction attacker (upper left corner), or provide no advantage to the attacker but perform worse than a majority classifier (lower right corner).
@@ 4.2 Results
Figure 4: Average training accuracy of $\varepsilon$-DP RFs with depth-5 trees as a function of the reconstruction error, for different privacy budgets $\varepsilon$ and forest sizes $\lvert\mathcal{T}\rvert$ on the COMPAS dataset
Result 4. Trade-offs between the RF predictive performance and the reconstruction error are complex.
Detailed trade-offs between the RFs accuracy and reconstruction error for all considered datasets and maximum depths are provided in Appendix C. Consistent with Figure 4, the only configurations yielding both non-trivial predictive performance and offering no advantage to the reconstruction adversary (upper right corner) are typically the smallest forests ($1$ tree) or the largest forests ($30$ trees). A plausible explanation is that small RFs inherently limit the amount of available information, providing natural protection against reconstruction. Conversely, in large RFs, the increased noise is offset by greater diversity among the trees, enhancing empirical privacy while preserving predictive performance. In contrast, medium-sized RFs often strike a poor balance: they provide enough diversity to aid reconstruction attacks but insufficient noise to effectively counteract them. These findings emphasize the importance of carefully selecting parameters such as forest size, tree depth, and privacy budget when training DP RFs for sensitive applications.
Result 5. Knowledge of the training set size $N$ is unnecessary.
@@
While prior work often assumes knowledge of the training set size $N$, Appendix A presents a modified version of our CP model that enables reconstruction without this information. Our experiments demonstrate that accurate reconstruction remains feasible under this assumption. Although the computational cost and reconstruction error increase slightly compared to the traditional setup where $N$ is known, the attack remains highly effective for most practically relevant $\varepsilon$ values.

### v2

@@ 4 Experimental Study
We now empirically evaluate our proposed reconstruction attack, and investigate whether it succeeds in accurately inferring a RF training data despite the DP protection. The source code and data needed to reproduce all our experiments and figures are available at https://github.com/vidalt/DRAFT-DP, under a MIT license.
@@ 4.1 Experimental Setup
Datasets and target models. We rely on three datasets, using the same data preprocessing as in Ferry et al. [2024b]. The COMPAS dataset [Angwin et al., 2016] contains records of $7{,}206$ criminal offenders from Florida, each characterized by $15$ binary attributes, and is used for recidivism prediction. The UCI Adult Income dataset [Dua and Graff, 2017] gathers data on $48{,}842$ individuals from the 1994 U.S. census. Each individual is described by $20$ binary features, with the objective of predicting whether an individual earns more than \$50K/year. Finally, the Default of Credit Card Clients (Default Credit) dataset [Yeh and hui Lien, 2009] includes $29{,}986$ customer records from Taiwan, with $22$ binary attributes, to predict payment defaults.
@@
The $\varepsilon$-DP RFs are trained as described in Section 2.2. In practice, we build upon the efficient and modular implementation of IBM’s diffprivlib library [Holohan et al., 2019], which we slightly adapt for our experiments. Following Ferry et al. [2024b], each experiment uses a randomly sampled training set of $N=100$ examples. While this sample size is in line with the literature and allows all experiments to be conducted in a moderate time frame, we also report results with larger training datasets in Appendix E, demonstrating the scalability of the approach. We build DP RFs of size $\lvert\mathcal{T}\rvert\in\{1,5,10,20,30\}$ and of maximum depth $d\in\{3,5,7\}$. Empirical evidence from Fan et al. [2003] suggests that the predictive accuracy of a random forest typically plateaus beyond $30$ trees, with $10$ to $30$ trees being sufficient for robust performance. This is particularly true in the DP setting, in which adding trees consumes additional privacy budget [Fletcher and Islam, 2019]. Finally, we evaluate privacy parameters $\varepsilon\in\{0.1,1,5,10,20,30\}$, ranging from very tight to large privacy budgets, which reflect commonly used values in practice [Desfontaines, 2021].
@@
Reconstruction attack. We use the OR-Tools CP-SAT solver [Perron and Didier, 2024] (v9) to solve the CP formulation introduced in Section 3. Each model resolution is limited to a maximum of $2$ hours of CPU time using $16$ threads with up to $7$ GB of RAM per thread. Note that in practice, feasible solutions – i.e., reconstructions that are compatible with the RF’s structure and the observed noisy counts – can be obtained using only a small fraction of the $2$-hours timeout (in average, less than two minutes), as reported in Appendix C.1. All experiments are run on a computing cluster over a set of homogeneous nodes with AMD EPYC 7532 (Zen 2) @ 2.40 GHz CPU.
@@
We consider two baselines in our experiments. As in Ferry et al. [2024b], a random reconstruction baseline is used to quantify the amount of information that can be inferred without knowledge of the RF (i.e., only based on the number of examples $N$, the different attributes $M$, their domains and one-hot encoding). This baseline randomly guesses the value of each attribute of each example while remaining consistent with one-hot encoding when applicable (we average its performance over $100$ runs). To quantify the effect of DP on the reconstruction error, we also run the state-of-the-art DRAFT attack against the trained RFs without the DP protection (i.e., with the true leaves’ counts).
@@
In line with the literature [Ferry et al., 2024b], while a reconstructed dataset is inferred, first it is aligned with the original training set using a minimum cost matching with the Manhattan distance, in order to identify which reconstructed examples correspond to which original ones. Then, the proportion of binary attributes that differ is averaged over the entire dataset to quantify the reconstruction error. Each experiment is repeated using five different random seeds and the results are averaged.
4.2 Results
Result 1. Reconstruction attacks succeed against DP RFs, even with meaningful privacy guarantees. Table 1 reports the reconstruction error for all our experiments. For commonly protective privacy budgets (e.g., $\varepsilon=1$ or $5$), our attack achieves significantly lower reconstruction error compared to the random baseline. This indicates that the adversary was able to gain insights from the RF to better infer the training data. Nonetheless, the reconstruction error consistently increases when reducing the privacy budget $\varepsilon$, which suggests that DP partially mitigates the reconstruction success. To assess this, Figure 3 compares the effectiveness of our reconstruction to that of DRAFT [Ferry et al., 2024b] applied on the same RF (without DP). For very tight privacy budgets (e.g., small $\varepsilon$), DP significantly increases the reconstruction error compared to the DRAFT baseline. However, as $\varepsilon$ increases, the protective effect of DP diminishes, and the difference in reconstruction error between our attack and DRAFT becomes much smaller, even for $\varepsilon$ values commonly considered protective.
Figure 2: Average reconstruction error as a function of the privacy budget $\varepsilon$ used to fit the target RF, for different numbers of depth-5 trees $|\mathcal{T}\rvert$ on the COMPAS dataset. For comparison, we also report the reconstruction error of DRAFT applied to the same RFs without DP protection.
Result 2. DP leads to a complex trade-off between the RFs’ size and the reconstruction attack success. As can be seen in Figure 3, larger DP RFs do not necessarily provide more information to the attacker, unlike non-DP RFs. A threshold effect is observed: the reconstruction error decreases as the number of trees increases, up to a point (typically around 5-10 trees), before increasing again. This behavior is consistent across all privacy budgets and is likely due to the privacy budget being distributed across more trees, resulting in noisier counts for individual leaves. Similarly, while deeper trees in the non-DP setting provide more information to the attacker, this is not always the case for DP RFs (see also, Table 1) since smaller leaves’ counts are proportionally more affected by the noise.
Result 3. Protection against reconstruction attacks comes at the expense of predictive accuracy. As shown in Table 1, the considered DP mechanism effectively mitigates the success of reconstruction attacks at very low $\varepsilon$ values (e.g., $\varepsilon=0.1$). However, at these privacy levels, the resulting RFs have poor predictive accuracy, often performing no better than a majority prediction classifier, as illustrated in Figure 3. Notably, only RFs with trivial predictive performance provide no advantage to the reconstruction adversary (compared to the random reconstruction baseline). This is also visible in Figure 4, in which most RFs either perform better than a majority classifier but provide a significant advantage to the reconstruction attacker (upper left corner), or provide no advantage to the attacker but perform worse than a majority classifier (lower right corner).
@@ 4.2 Results
Table 1: Average reconstruction error for different numbers of trees $\lvert\mathcal{T}\rvert$ , tree depths $d$ and privacy budgets $\varepsilon$ for the three datasets. Reconstruction errors worse than the random baseline are bold. Cases where the solver failed to find a feasible reconstruction within the time limit are marked with “–”.
@@
$\lvert\mathcal{T}\rvert=1$ $\varepsilon=0.1$ 0.26 0.24 0.22 0.29 0.28 0.21 0.35 0.31 0.29
$\varepsilon=1$ 0.25 0.20 0.17 0.26 0.20 0.14 0.32 0.29 0.24
$\varepsilon=5$ 0.24 0.23 0.30 0.25 0.20 0.18 0.32 0.31 0.25
$\varepsilon=10$ 0.24 0.22 0.33 0.25 0.21 0.18 0.32 0.30 0.27
$\varepsilon=20$ 0.24 0.19 0.27 0.25 0.17 0.20 0.32 0.28 0.30
$\varepsilon=30$ 0.24 0.19 0.27 0.25 0.16 0.18 0.32 0.29 0.31
$\lvert\mathcal{T}\rvert=5$ $\varepsilon=0.1$ 0.26 0.24 0.24 0.23 0.24 0.23 0.32 0.27 0.29
$\varepsilon=1$ 0.19 0.18 0.20 0.14 0.13 0.12 0.23 0.23 0.22
$\varepsilon=5$ 0.17 0.14 0.11 0.10 0.07 0.05 0.22 0.16 0.15
$\varepsilon=10$ 0.17 0.13 0.10 0.10 0.06 0.04 0.22 0.16 0.13
$\varepsilon=20$ 0.16 0.12 0.08 0.09 0.05 0.02 0.21 0.15 0.11
$\varepsilon=30$ 0.16 0.13 0.09 0.10 0.05 0.03 0.21 0.16 0.11
$\lvert\mathcal{T}\rvert=10$ $\varepsilon=0.1$ 0.28 0.31 0.34 0.25 0.28 0.28 0.32 0.31 0.30
$\varepsilon=1$ 0.17 0.19 0.23 0.12 0.14 0.16 0.21 0.24 0.24
$\varepsilon=5$ 0.15 0.12 0.15 0.07 0.07 0.06 0.15 0.16 0.19
$\varepsilon=10$ 0.13 0.10 0.11 0.06 0.05 0.04 0.14 0.13 0.13
$\varepsilon=20$ 0.14 0.09 0.09 0.06 0.03 0.03 0.14 0.12 0.11
$\varepsilon=30$ 0.13 0.08 0.09 0.05 0.03 0.02 0.14 0.11 0.09
$\lvert\mathcal{T}\rvert=20$ $\varepsilon=0.1$ 0.28 0.34 0.30 0.29 0.34 0.33 0.29 0.35 0.33
$\varepsilon=1$ 0.19 0.22 0.22 0.16 0.19 0.19 0.23 0.25 0.26
$\varepsilon=5$ 0.12 0.16 0.19 0.09 0.08 0.14 0.17 0.19 0.21
$\varepsilon=10$ 0.11 0.11 0.14 0.06 0.06 0.10 0.14 0.14 0.18
$\varepsilon=20$ 0.11 0.09 0.10 0.05 0.05 0.05 0.12 0.12 0.13
$\varepsilon=30$ 0.10 0.09 0.08 0.04 0.03 0.04 0.12 0.11 0.12
$\lvert\mathcal{T}\rvert=30$ $\varepsilon=0.1$ 0.34 0.35 0.33 0.33 0.36 0.38
$\varepsilon=1$ 0.21 0.22 0.25 0.17 0.20 0.21 0.23 0.26 0.25
$\varepsilon=5$ 0.15 0.16 0.22 0.09 0.10 0.16 0.17 0.20 0.24
$\varepsilon=10$ 0.13 0.14 0.18 0.07 0.08 0.13 0.14 0.14 0.19
$\varepsilon=20$ 0.10 0.10 0.13 0.06 0.06 0.09 0.13 0.13 0.17
$\varepsilon=30$ 0.10 0.09 0.11 0.04 0.04 0.06 0.11 0.11 0.13
@@
Figure 4: Average training accuracy of $\varepsilon$-DP RFs with depth-5 trees as a function of the reconstruction error, for different privacy budgets $\varepsilon$ and forest sizes $\lvert\mathcal{T}\rvert$ on the COMPAS dataset.
Result 4. For typical values of $\varepsilon$, one can infer information specific to the DP RF training set beyond general distributional patterns. The results in Table 1 demonstrate that the proposed attack extracts information useful for accurately reconstructing the DP RF training dataset $D{}$, which may reflect general distributional patterns and/or dataset-specific details. While the former is still valuable—especially given that our attack assumes no knowledge of the data distribution—privacy leakages are typically defined in terms of recovering information specific to the actual training set [Cohen et al., 2025]. Appendix D presents a complementary analysis in which we assess whether the reconstructed dataset is closer to $D{}$ than to other datasets from the same data distribution, in a statistically significant manner. The results suggest that for very tight DP budgets (i.e., $\varepsilon=0.1$), only distributional information is inferred while larger DP budgets (i.e., $\varepsilon\geq 5$) usually allow the retrieval of dataset-specific details. Between these two extrema, several configurations with small DP budgets (i.e., $\varepsilon=1$) can also lead to the inference of information specific to $D{}$. This emphasizes the need to thoroughly tune all hyperparameters beyond the sole DP budget.
@@
Result 5. Knowledge of the training set size $N$ is unnecessary. While prior work often assumes knowledge of the training set size $N$, Appendix A presents a modified version of our CP model that enables reconstruction without this information. Our experiments demonstrate that accurate reconstruction remains feasible under this assumption. Although the computational cost and reconstruction error increase slightly compared to the traditional setup in which $N$ is known, the attack remains highly effective for most practically relevant $\varepsilon$ values.
Result 6. The reconstruction attack scales well with the number of training examples $N$. In the Appendix E, we report results of additional experiments varying the size of the reconstructed training set $N$. They demonstrate that increasing the number of training examples $N$ improves the predictive performance and generalization of the built DP RFs, but also significantly benefits the reconstruction attack. For instance, the attack’s average reconstruction error against DP RFs with $\lvert\mathcal{T}\rvert=10$ depth-$5$ trees and $\varepsilon=5$ decreases from $0.16$ to $0.11$ when increasing the number of training examples from the Default of Credit Card Clients dataset from $N=100$ to $N=500$. This is due to the fact that the relative impact of the noise added to enforce DP decreases as $N$ increases, leading to more accurate leaf counts benefiting both the RF predictive performance and the proposed reconstruction attack.
Result 7. Knowledge of part of the training set attributes does not help reconstruct the others. In Appendix F, we report results of additional experiments assuming knowledge of part of the attributes, for all training examples. This setup corresponds to the situation in which part of the examples’ attributes are public information, and prior work on dataset reconstruction from (non-DP) RFs [Ferry et al., 2024b] demonstrated that knowing them facilitates the reconstruction of the others. We consistently observe that it is not the case when handling DP RFs: regardless of how many attributes the adversary knows, no advantage is gained to better reconstruct the remaining ones. Therefore, while DP does not entirely prevent the extraction of information useful for reconstructing the RF training dataset, it offers an interesting complementary protection by limiting the adversary’s ability to exploit auxiliary information.